Problem 4-26
Reaching a Financial Goal

You need to accumulate $10,000. To do so, you plan to make deposits of $1,200 per year - with the first payment being made a year from today - into a bank account that pays 7.07% annual interest. Your last deposit will be less than $1,200 if less is needed to round out to $10,000. How many years will it take you to reach your $10,000 goal? Round your answer up to the nearest whole number.
year(s)

How large will the last deposit be? Round your answer to the nearest cent.
$  
Answer #1

Using financial calculator
FV=10000
PMT=-1200
I/Y=7.07%
PV=0
CPT N=6.78073

It will take 7 years

Last deposit=10000-1200/7.07%*(1.0707^6-1)*1.0707=$792.94